# OpenFGL: A Comprehensive Benchmarks for Federated Graph Learning

### v1

@@ 3.2. Method-level FGL Algorithms
GNN Backbones. To provide a broader spectrum of graph learning paradigms on the client side, OpenFGL integrates a diverse range of local GNN backbones. Specifically, we implement various well-designed polling strategies (TopKPooling (Gao and Ji, 2019), SAGPooling (Lee et al., 2019), EdgePooling (Diehl et al., 2019), and PANPooling (Ma et al., 2020)) based on the most representative GIN (Xu et al., 2019) with weight-free MeanPooling (Xu et al., 2019) for Graph-FL. As for Subgraph-FL, OpenFGL includes prevalent GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017a), SGC (Wu et al., 2019) and GCNII (Chen et al., 2020b). The detailed descriptions of these backbones can be found in  (Ope, 2024) (A.2).
FL/FGL Algorithms. To achieve federated training, multi-client collaboration algorithms based on a trustworthy server are also crucial. Therefore, we follow the historical progression of FL to include a spectrum of algorithms from the most representative methods in CV to FGL: (1) CV-based FL algorithms: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), Scaffold (Karimireddy et al., 2020), MOON (Li et al., 2021), FedDC (Gao et al., 2022), FedProto (Tan et al., 2022), FedNH (Dai et al., 2023) and FedTGP (Zhang et al., 2024a); (2) FGL algorithms: GCFL+ (Xie et al., 2021) and FedStar (Tan et al., 2023) in Graph-FL and FedSage+ (Zhang et al., 2021), Fed-PUB (Baek et al., 2023), FedGTA (Li et al., 2023b), FGSSL (Huang et al., 2023), FedGL (Chen et al., 2024), AdaFGL (Li et al., 2024b), FGGP (Wan et al., 2024), FedDEP (Zhang et al., 2024b), and FedTAD (Zhu et al., 2024) in Subgraph-FL. More details about these algorithms can be found in  (Ope, 2024) (A.2). Notably, these algorithms are implemented with a unified API to facilitate future development in a user-friendly manner. For more details about our API design from the algorithm perspective, please refer to Sec. 4.5.
@@ 3.3. Experiment-level FGL Evaluations
Data Analysis. To delve into the data heterogeneity of FGL and offer essential insights for future studies, we advocate further analyzing the distributed graph data, encompassing the following aspects: (1) Feature KL Divergence: It reveals feature skew among clients while the label domain remains consistent, which may arise due to the different geographical locations of clients, such as the characteristics of a certain disease may significantly differ across various regions. (2) Label Distribution: As a more generalized data heterogeneity in CV-based FL, it is extensively discussed. However, in graph-based scenarios, the relationship between labels and topology frequently reveals underlying connections, characterized as homophily. Thus, we further integrate multi-level homophily metrics (Pei et al., 2020; Lim et al., 2021; Platonov et al., 2023b, a; Luan et al., 2022) to offer comprehensive analysis. (3) Topology Statics: This perspective stems from the critical role of topology in GNNs, especially in distributed scenarios. This arises from the significant impact of diverse local topology statistics on the local model, causing complex cascading effects in collaboration. Therefore, we examine topological differences among clients (e.g., Degree, Centrality, Largest Component) to provide insights.
@@
Robustness. To evaluate the practical deployment of FGL in real-world scenarios, we examine its robustness from the following perspectives: (1) Noise: This corresponds to data quality issues resulting from data collection. (2) Sparsity: This reflects scenarios with incomplete features, labels, and topology due to data scarcity and high labor costs, and with a low rate of client participation. (3) Client Communication: This simulates scenarios with network constraints or high communication costs. (4) Generalization: This evaluates generalizability of algorithms in intricate applications. (5) Privacy Preserve: This reflects the applicability of algorithms in privacy-sensitive scenarios, thereby we conduct an in-depth analysis from the perspective of Differential Privacy (DP). Please refer to  (Ope, 2024) (A.4) for further details on the robustness settings.
@@ 4.1. Performance Comparison
Graph-FL Scenario. Since the FGL research is still immature, the baselines for Graph-FL are scarce (only GCFL+ and FedStar). Therefore, we expand the number of local pooling-based backbones and CV-based FL algorithms, providing comprehensive results in Table 6. For Q1, we find that the benefits of federated collaboration are not significant for small-scale MUTAG, BZR, and COX2 due to their scarce data samples that can not support federated training and thus affect predictions. Subsequently, we conclude that C1: Federated collaboration is more advantageous for larger-scale datasets, benefiting from abundant data sources. As for Q2, we observe that: (1) GCFL+ and FedStar concentrate on topology Non-iid within the cross-domain simulation, thereby not consistently achieving competitive performance in this single-source setting. Therefore, they perform less favorably than FL algorithms on ENZYMES, COLLAB, and MULTI. (2) Existing FGL algorithms heavily rely on node semantics. We observe significant improvements on datasets with abundant node descriptions like DD and PROTEINS, whereas limited performance on BINARY whose node representations are initialized with node degrees. Consequently, we conclude that C2: Graph-FL algorithms still have improvement space, especially in single-source domain and constrained data semantics.
@@
Subgraph-FL Scenario. Compared to Graph-FL, Subgraph-FL has significantly advanced and offers numerous SOTA baselines. Their results are shown in Table 7. For Q1, although Subgraph-FL can benefit from abundant data samples, for Chameleon, Actor, and Ratings, the improvement from both FL and FGL algorithms is limited or even worse than solely local training in some cases. We attribute this to the well-known heterophily challenges, where differences in node connection rules across clients significantly affect local updates and server-side collaboration, deviating from the global optimum and resulting in sub-optimal predictive performance. Although AdaFGL addresses this issue through personalized technology, there is still room for better performance. Based on this and C1, we conclude that C3: The prerequisite for positive impacts of federated collaboration is the uniform distribution of node features, labels, and topology. Regarding Q2, we observe that: (1) Subgraph-FL is thriving, with numerous baselines vigorously competing for the best performance. Among them, FedTAD and AdaFGL stand out in most cases. (2) The outstanding performance of existing algorithms stems from the fine-grained exploration of the topology, but some methods lack scalability when dealing with large-scale ogbn-products, resulting in OOM (out-of-memory) errors. Consequently, we conclude that C4: Subgraph-FL algorithms should address the complexity of real-world deployment, especially focusing on large-scale scenarios and graph-specific federated heterogeneity challenges.
@@ 4.2. Robustness Analysis
Local Noise. To answer Q3 from the perspective of noise, the experimental results are shown in the upper part of Fig. 1(a)-(c). For feature noise, we randomly select the channels of node features and inject Gaussian noise. To simulate edge noise in the Graph-FL without node labels, we utilize Metattack (Zügner and Günnemann, 2019) to add noise edges, significantly perturbing the training gradients. As for label noise, we randomly reassign non-real labels to a certain proportion of training node samples. Based on the experimental results, we observe that FGL algorithms are highly sensitive to edge noise compared to topology-agnostic FL algorithms. This inherent limitation directly disrupts the model optimization of GCFL+ and FedStar, misleading local updates and topology-driven collaboration, thereby resulting in sub-optimal predictive performance. However, GCFL+ and FedStar demonstrate superior robustness under feature and label noise, as they address client interference with each other using server-side clustering customization and additional local models maintained at client-side. Consequently, we can conclude that C5: Noise scenarios determine the performance lower bound for FGL algorithms, where personalized strategies emerge as crucial technologies. However, they fall slightly short in addressing edge noise.
Local Sparsity. To answer Q3 from the perspective of sparsity, the experimental results are shown in the lower part of Fig. 1(a)-(c). Regarding feature sparsity, we simulate partial feature absence for unlabeled nodes and randomly remove the original edges. As for label sparsity, we change the ratio of training nodes. Under feature sparsity, FGSSL and Fed-PUB exhibit significant performance fluctuations due to heavy reliance on high-quality features for node-wise contrastive learning and model-wise gradient matching. Conversely, FedSage+, AdaFGL, and FedTAD leverage multi-client collaboration (i.e., subgraph generation, augmentation, and knowledge distillation), mitigating confusion in under-trained model collaboration, thus ensuring robustness. Similar analysis can extend to label sparsity, as both of them directly affect local updates. Regarding edge sparsity, FedTAD’s vulnerability lies in its reliance on client-side topology embeddings to supervise server-sider pseudo-subgraph generator, rendering it fragile in extreme edge sparsity scenarios. However, AdaFGL and FedSage+, leveraging topology mining, adeptly handle this challenge. Based on this, we can conclude that: C6: Sparsity scenarios determine the performance upper bound for FGL algorithms, where multi-client collaboration is the pivotal technology, particularly in synergy with topology mining.
Client Participation. To answer Q4, we present the experimental results in Fig. 1(d), where robust FGL algorithms with low client participation exhibit one of the following characteristics: (1) They rely less on messages received from the server and focus on well-designed local training mechanisms; (2) They custom global messages for each participating client for local personalized training.
For instance, in Graph-FL, the unstable performance of FedStar arises from its heavy reliance on specific topological properties and well-learned structure embeddings. Conversely, GCFL+ ensures high-quality local updates by tailoring the most suitable messages for each client through server-side client clustering. In Subgraph-FL, AdaFGL, Fed-PUB, and FedGTA rely on client-side personalized local training and server-side pseudo graph-driven trainable weight clustering or identification of subgraph statistics to ensure custom global messages for each participating client. Consequently, we can conclude that C7: Low client participation underscores the emphasis of FGL algorithms on local updates, highlighting the importance of local data understanding and customizing messages for each client.
@@
Generalization. To answer Q5, the experimental results are shown in Table 8, where graph-level cross-domain setting includes MUTAG, COX2, PTC-MR, AIDS, ENZYMES, DD, PROTEINS, COLLAB, and IMDB-B/M and subgraph-level setting includes Cora, CiteSeer, Pubmed, CS, and Physics. We observe that current FGL algorithms exhibit inconsistent performance across data simulations. Specifically, in Graph-FL, tiny-scale datasets with abundant node descriptions such as DHFR and DD mitigate inherent differences among data simulations potentially due to over-fitting issues. As for the Subgraph-FL, data simulations present challenges for existing FGL algorithms. For instance, Fed-PUB and AdaFGL, incorporating personalized strategies, lose their previous advantages, whereas FedSage+ and FedGTA, emphasizing multi-client collaboration mechanisms, show significant potential. Therefore, we can conclude that C8: In practical deployments aiming for generalization, client-specific design should be used cautiously, with an emphasis on discovering inherent global consensus.
@@
DP-based Privacy Preserve. To answer Q6, the experimental results are shown in Table 9, where $\epsilon$ is the privacy budget. To implement DP in FGL, we introduce well-selected random noise in the local model gradients for server-side perturbed model aggregation. More technology details can be found in (Ope, 2024) (A.8). Meanwhile, we integrate FedDEP into OpenFGL, which introduces additional edge-level DP. Based on the results, we observe that a large $\epsilon$ enables privacy-preserving methods to match the accuracy of the non-private approach. However, reducing $\epsilon$ results in a notable performance drop, primarily due to the need for local models for excessive noise injection into gradients, leading to significant degradation. Regarding FedDEP, compared to FedSage+, it achieves edge-level differential privacy through random sampling and enhances model capacity with a deep neighbor generation module, striking a balance between performance and privacy preservation. Based on this, we conclude that C9: FGL algorithms currently face a dilemma between predictive performance and privacy preservation.
@@ 4.3. FGL Algorithm Complexity Analysis
To progressively illustrate the theoretical algorithm complexity analysis, we choose SGC (Wu et al., 2019) as the local model ($k$-step feature propagation), otherwise, we adopt the model architecture ($L$-layer) used in their original paper. For the $k$-layer SGC model with batch size $b$, the $\mathbf{X}^{(k)}$ is the propagated feature bounded by a space complexity of $O((b+k)f)$. The overhead for linear regression by multiplying $\mathbf{W}$ is $O(f^{2})$. In the training stage, the above procedure is repeated to iteratively update the model weights. For the server performing FedAvg, it needs to receive the model weights and the number of samples participating in this round. Its space complexity and time complexity are bounded as $O(Nf^{2})$ and $O(N)$. As discovered by previous studies (Chen et al., 2020a; Li et al., 2024a; Zhang et al., 2022), the dominating term is $O(kmf)$ or $O(Lmf)$ when the graph is large since feature learning can be accelerated by parallel computation. Consequently, $O(Lmf)$ emerges as the dominating complexity term of linear transformation, and the execution of full neighbor propagation becomes the primary bottleneck for achieving scalability. Although FGL offers a new perspective for large-scale graph learning through a distributed paradigm, it still requires the deployment of suitable scalable GNNs in billion-level scenarios on the local client.

### v2

@@ 3.2. Method-level FGL Algorithms
GNN Backbones. To provide a broader spectrum of learning paradigms on the client side, OpenFGL integrates a diverse range of local GNN backbones. Specifically, we implement various well-designed polling strategies (TopKPooling (Gao and Ji, 2019), SAGPooling (Lee et al., 2019), EdgePooling (Diehl et al., 2019), and PANPooling (Ma et al., 2020)) based on the most representative GIN (Xu et al., 2019) with weight-free MeanPooling (Xu et al., 2019) for Graph-FL. As for Subgraph-FL, OpenFGL includes prevalent GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017a), SGC (Wu et al., 2019) and GCNII (Chen et al., 2020b). The detailed descriptions of these backbones can be found in  (Ope, 2024) (A.2).
FL/FGL Algorithms. To achieve federated training, multi-client collaboration algorithms are crucial (CV-based FL also can be applied to FGL). We follow the historical progression of FL to include a spectrum of algorithms from the most representative methods in CV to FGL: (1) CV-based FL algorithms: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020b), Scaffold (Karimireddy et al., 2020), MOON (Li et al., 2021a), FedDC (Gao et al., 2022), FedProto (Tan et al., 2022), FedNH (Dai et al., 2023) and FedTGP (Zhang et al., 2024b); (2) all FGL algorithms possible: GCFL+ (Xie et al., 2021) and FedStar (Tan et al., 2023) in Graph-FL and FedSage+ (Zhang et al., 2021), Fed-PUB (Baek et al., 2023), FedGTA (Li et al., 2023b), FGSSL (Huang et al., 2023), FedGL (Chen et al., 2024), AdaFGL (Li et al., 2024e), FGGP (Wan et al., 2024), FedDEP (Zhang et al., 2024c), and FedTAD (Zhu et al., 2024) in Subgraph-FL. More details about these algorithms can be found in  (Ope, 2024) (A.2). Notably, these algorithms are implemented with a unified API to facilitate future development in a user-friendly manner. For more details about our API design from the algorithm perspective, please refer to Sec. 4.5.
@@ 3.3. Experiment-level FGL Evaluations
Data Analysis. (1) Feature KL Divergence: It reveals feature skew among clients while the label domain remains consistent, which may arise due to the different geographical locations of clients, such as the characteristics of a certain disease may significantly differ across various regions. (2) Label Distribution: It is widely discussed in CV-based FL. However, in FGL, the relationship between labels and topology frequently reveals underlying connections, characterized as homophily. Thus, we further integrate multi-level homophily metrics (Pei et al., 2020; Lim et al., 2021; Platonov et al., 2023b, a; Luan et al., 2022) to offer comprehensive analysis. (3) Topology Statics: This perspective stems from the critical role of topology in GNNs, especially in distributed scenarios. This arises from the significant impact of diverse local topology statistics on the local model, causing complex cascading effects in collaboration. Therefore, we examine topological differences among clients (e.g., Degree, Centrality, Largest Component) to provide insights.
@@
Robustness. To evaluate the practical deployment of FGL, we examine its robustness from the following perspectives: (1) Noise: This corresponds to data quality issues resulting from data collection (Li et al., 2020a; Liu et al., 2022). (2) Sparsity: This reflects scenarios with incomplete features, labels, and topology due to data scarcity and high labor costs, and with a low rate of client participation (Gu et al., 2023a; Wu et al., 2024b; Li et al., 2024a). (3) Client Communication: This simulates scenarios with network constraints or high communication costs (Yao et al., 2023; Yuan et al., 2024). (4) Generalization: This evaluates the effectiveness of algorithms in various scenarios. (5) Privacy Preserve: This reflects the applicability of algorithms in privacy-sensitive scenarios, thereby we conduct an in-depth analysis from the perspective of Differential Privacy (DP). Please refer to  (Ope, 2024) (A.4) for further details on the robustness settings.
@@ 4.1. Performance Comparison
Graph-FL Scenario. Since the baselines for Graph-FL are scarce. Therefore, we expand pooling-based backbones and CV-based FL in Table 6. For Q1, we find that the benefits of federated collaboration are not significant for small-scale MUTAG, BZR, and COX2 due to limited data that can not support federated training and thus affect predictions. Subsequently, we conclude that C1: Federated collaboration is more advantageous for larger-scale datasets, benefiting from abundant data sources (Ruan et al., 2024; Li et al., 2024b). As for Q2, we observe that: (1) GCFL+ and FedStar concentrate on topology Non-iid within the cross-domain simulation, thereby not consistently achieving competitive performance in this single-source setting. Therefore, they perform less favorably than FL algorithms on ENZYMES, COLLAB, and MULTI. (2) Existing FGL algorithms heavily rely on node semantics. We observe significant improvements on datasets with abundant node descriptions like DD and PROTEINS, whereas limited performance on BINARY whose node representations are initialized with node degrees. Consequently, we conclude that C2: Graph-FL algorithms still have improvement space, especially in the single-source domain and constrained data semantics (Abushofa et al., 2023; Li et al., 2024d).
@@
Subgraph-FL Scenario. Experimental results are shown in Table 7. For Q1, although Subgraph-FL can benefit from abundant data, for Chameleon, Actor, and Ratings, the improvement from both FL and FGL is limited or even worse than solely local training in some cases. We attribute this to the heterophily, where differences in node connection rules across clients significantly affect local updates and server-side collaboration, deviating from the global optimum and resulting in sub-optimal performance. Although AdaFGL addresses this issue through personalized technology, there is still room for better performance. Based on this and C1, we conclude that C3: The prerequisite for positive impacts of federated collaboration is the uniform distribution of node features, labels, and topology (Liu et al., 2024; Xia et al., 2024). Regarding Q2, we observe that: (1) Subgraph-FL is thriving, with numerous baselines vigorously competing for the best performance. Among them, FedTAD and AdaFGL stand out in most cases. (2) The outstanding performance of existing algorithms stems from the fine-grained exploration of the topology, but some methods lack scalability when dealing with large-scale ogbn-products, resulting in OOM (out-of-memory) errors. Consequently, we conclude that C4: Subgraph-FL algorithms need to resolve the complexity in real-world deployments, especially focusing on large-scale scenarios and graph-specific federated heterogeneity challenges (Li et al., 2023c, 2024f).
@@ 4.2. Robustness Analysis
Local Noise. To answer Q3 from the perspective of noise, the experimental results are shown in the upper part of Fig. 1(a)-(c). For feature noise, we randomly select the channels of node features and inject Gaussian noise. To simulate edge noise in the Graph-FL without node labels, we utilize Metattack (Zügner and Günnemann, 2019) to add noise edges, significantly perturbing the training gradients. As for label noise, we randomly reassign non-real labels to a certain proportion of training node samples. Based on the experimental results, we observe that FGL algorithms are highly sensitive to edge noise compared to topology-agnostic FL algorithms. This inherent limitation directly disrupts the model optimization of GCFL+ and FedStar, misleading local updates and topology-driven collaboration, thereby resulting in sub-optimal predictive performance. However, GCFL+ and FedStar demonstrate superior robustness under feature and label noise, as they address client interference with each other using server-side clustering customization and additional local models maintained at client-side. Consequently, we can conclude that C5: Noise scenarios determine the performance lower bound for FGL algorithms, where personalized strategies emerge as crucial technologies. However, they fall slightly short in addressing edge noise (Zhang et al., 2024a; Lin et al., 2024).
Local Sparsity. To answer Q3 from the perspective of sparsity, the experimental results are shown in the lower part of Fig. 1(a)-(c). Regarding feature sparsity, we simulate partial feature absence for unlabeled nodes and randomly remove the original edges. As for label sparsity, we change the ratio of training nodes. Under feature sparsity, FGSSL and Fed-PUB exhibit significant performance fluctuations due to heavy reliance on high-quality features for node-wise contrastive learning and model-wise gradient matching. Conversely, FedSage+, AdaFGL, and FedTAD leverage multi-client collaboration, mitigating confusion in under-trained model collaboration, thus ensuring robustness. Similar analysis can extend to label sparsity, as both of them directly affect local updates. Regarding edge sparsity, FedTAD’s vulnerability lies in its reliance on client-side topology embeddings to supervise server-sider pseudo-subgraph generator. However, AdaFGL and FedSage+, leveraging topology mining, adeptly handle this challenge. Based on this, we can conclude that: C6: Sparsity scenarios determine the performance upper bound for FGL algorithms, where multi-client collaboration is the pivotal technology, particularly in synergy with topology mining (Wu et al., 2024a; Babakniya et al., 2023).
Client Participation. To answer Q4, we present the experimental results in Fig. 1(d), where robust FGL algorithms with low client participation exhibit one of the following characteristics: (1) They rely less on messages received from the server and focus on local training; (2) They custom global messages for each participating client. For instance, in Graph-FL, the unstable performance of FedStar arises from its heavy reliance on specific topological properties. Conversely, GCFL+ ensures high-quality local updates by tailoring the most suitable messages for each client through server-side clustering. In Subgraph-FL, AdaFGL, Fed-PUB, and FedGTA rely on client-side personalized training, server-side pseudo-graph-driven clustering, or identification of subgraph statistics to ensure custom global messages for each participating client. Consequently, we can conclude that C7: Low client participation underscores the emphasis of FGL algorithms on local updates, highlighting the importance of local data understanding and customizing messages for each client (Fu et al., 2023; Qu et al., 2023a).
Generalization. To answer Q5, the experimental results are shown in Table 8, where graph-level cross-domain setting includes MUTAG, COX2, PTC-MR, AIDS, ENZYMES, DD, PROTEINS, COLLAB, and IMDB-B/M and subgraph-level setting includes Cora, CiteSeer, Pubmed, CS, and Physics. We observe that current FGL algorithms exhibit inconsistent performance across data simulations. Specifically, in Graph-FL, tiny-scale datasets with abundant node descriptions such as DHFR and DD mitigate inherent differences among data simulations potentially due to over-fitting issues. As for the Subgraph-FL, data simulations present challenges for existing FGL algorithms. For instance, Fed-PUB and AdaFGL, incorporating personalized strategies, lose their previous advantages, whereas FedSage+ and FedGTA, emphasizing multi-client collaboration mechanisms, show significant potential. Therefore, we can conclude that C8: In practical deployments aiming for generalization, client-specific design should be used cautiously, with an emphasis on discovering inherent global consensus (Guo et al., 2023; Mora et al., 2024).
@@
DP-based Privacy Preserve. To answer Q6, the experimental results are shown in Table 9, where $\epsilon$ is the privacy budget. To implement DP in FGL, we introduce well-selected random noise in the local model gradients for server-side perturbed model aggregation. More technology details can be found in (Ope, 2024) (A.8). Meanwhile, we integrate FedDEP into OpenFGL, which introduces additional edge-level DP. Based on the results, we observe that a large $\epsilon$ enables privacy-preserving methods to match the accuracy of the non-private approach. However, reducing $\epsilon$ results in a notable performance drop, primarily due to the need for local models for excessive noise injection into gradients, leading to significant degradation. Regarding FedDEP, compared to FedSage+, it achieves edge-level differential privacy through random sampling and enhances model capacity with a deep neighbor generation module, striking a balance between performance and privacy preservation. Based on this, we conclude that C9: FGL algorithms currently face a dilemma between predictive performance and privacy preservation (Fu et al., 2024; Zhang et al., 2024f).
@@ 4.3. FGL Algorithm Complexity Analysis
Table 11. Performance (%) on large-scale GraphFL.
Graph-FL
Label (0.5)
molhiv
(ROC-AUC)
molpcba
ppa
code2
(F1 score)
FedAvg 72.2±2.4 20.4±0.4 65.8±1.6 29.8±0.4
FedProx 72.8±2.2 20.9±0.3 65.4±1.9 29.5±0.5
Scaffold 71.7±2.5 21.2±0.3 66.4±1.5 29.7±0.5
FedProto 70.5±1.8 19.3±0.2 63.7±1.2 28.6±0.3
FedNH 71.2±2.0 19.5±0.2 64.4±1.1 28.4±0.2
FedTGP 70.8±2.4 19.7±0.3 64.2±1.4 28.8±0.3
GCFL+ 73.4±1.9 21.8±0.3 67.0±1.5 30.5±0.3
FedStar 73.8±1.7 21.5±0.2 67.2±1.8 31.2±0.4
Table 12. Efficiency on Louvain-based Physics (Subgraph-FL).
Method Test Acc (%) Memory Com. Running Time
Scaffold 90.6±0.5 538k 1076k 152.42s
FedTGP 87.4±0.3 539k 0.38k 76.87s
FedSage+ 91.7±0.8 1296k 1784k 1517.96s
GCFL+ 91.0±0.3 538k 538k 184.24s
Fed-PUB 91.6±0.5 1076k 1076k 391.60s
FedGTA 91.5±0.3 538k 539k 120.45s
FedStar 91.3±0.5 1076k 539k 268.12s
FGSSL 91.1±0.6 751k 540k 476.58s
AdaFGL 92.0±0.4 964k 538k 162.19s
@@
For convenience, we choose SGC (Wu et al., 2019) as the local model ($k$-step feature propagation), otherwise, we adopt the model architecture ($L$-layer) used in their original paper. For the $k$-layer SGC model with batch size $b$, the $\mathbf{X}^{(k)}$ is the propagated feature bounded by a space complexity of $O((b+k)f)$. The overhead for linear regression by multiplying $\mathbf{W}$ is $O(f^{2})$. In the training stage, the above procedure is repeated to iteratively update the model weights. For the server performing FedAvg, it needs to receive the model weights and the number of samples participating in this round. Its space complexity and time complexity are bounded as $O(Nf^{2})$ and $O(N)$. As discovered by previous studies (Chen et al., 2020a; Li et al., 2024c; Zhang et al., 2022), the dominating term is $O(kmf)$ or $O(Lmf)$ when the graph is large since feature learning can be accelerated by parallel computation. Consequently, $O(Lmf)$ emerges as the dominating complexity term of linear transformation. Although FGL offers a new perspective for large-scale graph learning through a distributed paradigm, it still requires the deployment of suitable scalable GNNs on the local client.
@@
Table 13. A summary and selection suggestions for current prevalent FGL studies.
Methods Effectiveness Robustness Efficiency
Statistic Heterogeneity Noise Sparsity Low Client Generalization Privacy Communication Scalability Parallelism
[…]